# Frequentist Validity of Epistemic Uncertainty Estimators

### v1

@@ Abstract
Decomposing prediction uncertainty into its aleatoric (irreducible) and epistemic (reducible) components is critical for the development and deployment of machine learning systems. A popular, principled measure for epistemic uncertainty is the mutual information between the response variable and model parameters. However, evaluating this measure requires access to the posterior distribution of the model parameters, which is challenging to compute. In view of this, we introduce a frequentist measure of epistemic uncertainty based on the bootstrap. Our main theoretical contribution is a novel asymptotic expansion that reveals that our proposed (frequentist) measure and the (Bayesian) mutual information are asymptotically equivalent. This provides frequentist interpretations to mutual information and new computational strategies for approximating it. Moreover, we link our proposed approach to the widely-used heuristic approach of deep ensembles, giving added perspective on their practical success.
@@ 1 Introduction
While machine learning systems are increasingly accurate in tightly-controlled settings, a key challenge to their deployment in more complex, real-world environments is uncertainty awareness—algorithms should be able to accurately report on their level of confidence or reliability in different conditions. Moreover, there are different underlying reasons for uncertainty. Uncertainty may be due to inherent unpredictability in a population with the current feature set (aleatoric or irreducible uncertainty) or instead due to limited training data (epistemic or reducible uncertainty). Distinguishing these types of uncertainty is important in order to take the correct downstream action. High aleatoric uncertainty indicates that the current prediction task is inherently challenging, and we cannot have high confidence in any one prediction unless we are able to collect richer data (not just more data). On the other hand, high epistemic uncertainty signals that collecting more training data in this region is likely to improve the model, which makes it suitable for guiding model development and active learning (Houlsby et al.,, 2011; Gal et al.,, 2017; Smith et al.,, 2023). We focus on this quantity in this work.
@@
A popular measure of epistemic uncertainty is the mutual information (MI) between model parameters and predictions which quantifies the expected reduction in uncertainty from observing new data (Houlsby et al.,, 2011; Depeweg et al.,, 2018). MI is a fundamentally Bayesian quantity that requires access to the posterior distribution over model parameters, which generally renders its exact calculation intractable. As such, substantial effort has gone into developing approximate Bayesian computational techniques (MacKay,, 1995; Graves,, 2011; Welling and Teh,, 2011; Hernandez-Lobato and Adams,, 2015; Gal and Ghahramani,, 2016), but they often require changes to model architectures or training procedures, which may compromise predictive accuracy.
As an alternative approach, our work proposes an estimator of MI based on the bootstrap. In particular, we provide an asymptotic expansion for MI which links it to the Fisher information and shows how epistemic uncertainty differs from frequentist variance. We use this to develop a bootstrap estimator which targets the MI. The estimator is simple—we form bootstrap samples and use them as if they were samples from the posterior distribution, computing the MI accordingly. This requires minimal engineering effort, and does not require any restrictions on the underlying architecture, making it an attractive complement to existing MI estimators. We conduct experiments to showcase the validity of the bootstrapping approach in a deep learning setting and the success of our measure on an active learning task.
Morever, our proposal allows for a natural decomposition of epistemic uncertainty into components arising due to data sampling, and due to stochasticity in the training procedure. This allows us to reason that deep ensembles capture the portion of epistemic uncertainty due to stochastic optimization, which empirically appears to be the majority component, explaining their success. Finally, we provide a proof-of-concept experiment showing how data attribution approaches can be leveraged to provide estimates of epistemic uncertainty.
@@ 1.1 Our Contribution
Our contribution can be summarized as follows.
@@
  • Bootstrap estimation of mutual information: We propose an estimator of MI based on bootstrap. This allows for study between frequentist notions of uncertainty which are based on variance under repeated draws of data, and Bayesian notions which are based on spread in distribution of the parameters.
  • Connection with deep ensembles: We propose a decomposition of our epistemic uncertainty estimator into components arising from training stochasticity and data sampling, and use this to provide insight into the success of deep ensembles.
@@ 1.2 Illustration
Figure 1: Teaser Simulation: (left) Training data, true label probability and posterior sample predictions. (center) Predictions from MLEs over bootstrapped datasets. (right) Epistemic uncertainty calculated through Bayesian inference (MCMC) versus our bootstrap based measure.
To provide the key intuition for our method, we provide the following simulation. Consider a binary logistic regression model, $\mathbb{P}(y=1|x,\theta)=1/(1+e^{-\theta_{1}x-\theta_{0}})$. We sample the features $x$ from a standard normal and also impose standard normal priors on $\theta_{0},\theta_{1}$. Fig. 1(left) visualizes the data and the true model along with 100 posterior samples from the true Bayesian posterior, obtained using a Markov chain Monte Carlo (MCMC) algorithm. Next, we calculate maximum likelihood estimates (MLEs) for the parameters across 100 bootstrapped datasets and show these in Fig. 1(middle)—note the remarkable similarity in the posterior and bootstrap samples. Finally, we use these MCMC and bootstrap samples to form our estimate of the epistemic uncertainty as per Algorithm 1 noting the excellent agreement in Fig. 1(right). In the remainder of this work, we develop this connection.
@@ 2.1 Mutual Information as Epistemic Uncertainty
A popular uncertainty decomposition (MacKay,, 1992; Houlsby et al.,, 2011; Depeweg et al.,, 2018) leverages information-theoretic quantities to propose measures of aleatoric and epistemic uncertainty. Let the random variables $Y$ and $X$ denote the labels and features respectively in a supervised learning setup. Let $\mathcal{D}_{n}$ be the data comprising of $n$ realizations of $(Y,X)$. Further, let $\theta$ be the random variable denoting the parameters of a prediction model. Information theory states that $\mathrm{H}(Y)$ is the uncertainty in $Y$ and defines mutual information as
@@
The quantification of epistemic uncertainty as the mutual information (MI) between the (test) data and parameters is conceptually pleasing since Eq. 1 reveals it to be the reduction in label uncertainty from knowing the model parameters. This MI is a fundamentally Bayesian object since it requires $\theta$ to be random. In our paper, we will treat the (Bayesian) MI as the targeted measure of epistemic uncertainty and shall investigate various estimators for it.
@@ 2.2 Bootstrap
The original bootstrap procedure (Efron,, 1979) was proposed to estimate the variance of an estimator under different draws of the data. Consider a setting where there are no repeated data points. Bootstrap first samples weights from a symmetric multinomial distribution with all event probabilities $1/n$. These weights are used to reweight the original data to create a “bootstrapped dataset” over which the estimator is recomputed. This procedure is replicated $B$ times to obtain $B$ “bootstrapped estimates”. The key intuition behind the bootstrap—the “plug-in” principle—proposes that the distribution of these bootstrapped estimates approximates the distribution of the original estimator under redraws of the data. As such, we can obtain an estimate of the variance of our estimator as the empirical variance of the bootstrapped estimates.
We can also instead draw weights from a symmetric Dirichlet distribution (Rubin,, 1981). We hypothesize that in a deep learning setting, these weights can be more effective at creating bootstrapped datasets since it allows the neural network to still see all the training instances and hence achieve better predictive performance than multinomial weights which omit training points entirely. For our paper, we thus use Dirichlet weights but emphasize that our theorems hold for multinomial weights too (Appendix A).
@@ 2.3 Deep Ensembles
Deep ensembles (Lakshminarayanan et al.,, 2017) are a simple but powerful uncertainty quantification technique for deep neural networks (DNNs). A deep ensemble is obtained by training multiple models with the exact same architecture and training data but with different random seeds controlling stochasticity in the training procedure (e.g. initial weights, data shuffling etc.). The deep ensemble model prediction is simply the average of predictions of member models. They have shown tremendous success at a variety of uncertainty quantification tasks (Ashukha et al.,, 2020; Ovadia et al.,, 2019).
@@ 3.1 Problem Setup
$\mathrm{I}\left(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n}\right)=\mathrm{H}\left(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n}\right)-\mathrm{H}\left(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n}\right)=\mathrm{H}\left(\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\hat{p}(X_{\mathrm{test}};\theta)\right]\right)-\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))\right].$
@@ 3.2 Asymptotics of Mutual Information
Theorem 1.
Under the assumptions needed for Bernstein von-Mises to hold and uniform integrability of the predicted probabilities ((A1)-(A7) in Appendix A) we have the asymptotic expansion
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=\frac{1}{2n}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}+o_{p}(n^{-1})\qquad\forall\ X_{\mathrm{test}},$ (2)
where $\sigma_{k}^{2}$ is
@@
Proof Sketch: Our analysis proceeds by Taylor expanding the entropy function about the mean posterior predicted probability. We then apply the functional delta method to the Bernstein von-Mises theorem to obtain the limiting law of our predictions. Recall that the Bernstein von-Mises theorem holds under correct model specification, sufficient smoothness, identifiability, and feasibility of $\theta_{0}$ under the prior (Theorem 10.1 of van der Vaart, (2000)). Finally, we use uniform integrability to obtain the limiting posterior moments. We defer the full proof to Appendix A.
Firstly, the theorem shows that mutual information decreases linearly in the number of samples. Next, it connects mutual information to the Fisher information and perhaps surprisingly shows how the first order asymptotics are governed by the inverse Fisher information divided by the true conditional probabilities.
@@
as noted in MacKay, (1992). Recall that $n^{-1}\mathcal{I}^{-1}(\theta_{0})$ is the (asymptotic) variance of the MLE. The numerator in the expression above can then be interpreted as the variance of the model prediction at $X_{\mathrm{test}}$. 1/2p(1p)1/2p(1-p) achieves a minima at $1/2$ and tends to infinity at 0 and $1$. The first order term is then directly proportional to the model variance but also inversely proportional to the the ground truth certainty (i.e. how close $\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})$ is to 0 or 1). Thus, we see that this first order term matches the intuition behind epistemic uncertainty; if our predictions are uncertain but the true outcome is certain then we can expect to reduce our uncertainty and hence have high epistemic uncertainty.
3.3 Our Bootstrap Estimator
Theorem 1 suggests that asymptotically correct estimation of the Fisher information and true model probabilities can thus be used to construct first order accurate estimators of mutual information. In light of this, we propose a bootstrap based estimator which we denote by $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$.
Our estimation procedure proceeds by first sampling weights $\xi$ from the symmetric Dirichlet with concentration parameter 1. Denote our MLE of $\theta$ over the bootstrapped dataset as $\hat{\theta}_{b}$ (for example, as obtained by minimizing cross-entropy loss). We denote the distribution of this estimate (induced by the randomness in the weights) given the data $\mathcal{D}_{n}$ as $p(\hat{\theta}_{b}|\mathcal{D}_{n})$. Then,
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n}):=\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
Practical calculation is incredibly easy as illustrated by Algorithm 1.
@@ 3.3 Our Bootstrap Estimator
Theorem 2.
Under the assumptions needed for Bernstein von-Mises to hold, for asymptotic normality of the MLE to hold and uniform integrability of the predicted probabilities ((A1)-(A12) in Appendix A), we have
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}}.$
Proof Sketch: We proceed as in Theorem 1 but invoke bootstrap limit laws (Theorem 10.16 of Kosorok, (2007)) instead of Bernstein von-Mises to obtain an identical asymptotic expansion for our estimator. We provide the complete proof in Appendix A.
@@
Theorem 2 proves the asymptotic validity of our estimator and justifies its use. Our procedure is model agnostic and can be used even when Bayesian approximations are ad-hoc or entirely unavailable. It thus provides a theoretically justified path to approximating Bayesian information-theoretic quantities using a well studied frequentist procedure.

### v2

# Deep Ensembles for Epistemic Uncertainty: A Frequentist Perspective

@@ Abstract
Decomposing prediction uncertainty into aleatoric (irreducible) and epistemic (reducible) components is critical for the reliable deployment of machine learning systems. While the mutual information between the response variable and model parameters is a principled measure for epistemic uncertainty, it requires access to the parameter posterior, which is computationally challenging to approximate. Consequently, practitioners often rely on probabilistic predictions from deep ensembles to quantify uncertainty, which have demonstrated strong empirical performance. However, a theoretical understanding of their success from a frequentist perspective remains limited. We address this gap by first considering a bootstrap-based estimator for epistemic uncertainty, which we prove is asymptotically correct. Next, we connect deep ensembles to the bootstrap estimator by decomposing it into data variability and training stochasticity; specifically, we show that deep ensembles capture the training stochasticity component. Through empirical studies, we show that this stochasticity component constitutes the majority of epistemic uncertainty, thereby explaining the effectiveness of deep ensembles.
Machine Learning, Uncertainty Quantification, Deep Learning, Epistemic Uncertainty, Deep Ensembles, Bootstrap, Mutual Information, Bayesian Frequentist Equivalences
@@ 1 Introduction
Figure 1: Teaser Simulation: (left) Training data, true label probability and posterior sample predictions. (center) Predictions from MLEs over bootstrapped datasets. (right) Epistemic uncertainty calculated through Bayesian inference (MCMC) versus the bootstrap estimate.
@@
While machine learning systems are increasingly accurate in tightly-controlled settings, a key challenge to their deployment in more complex, real-world environments is uncertainty awareness—algorithms should be able to accurately report on their level of confidence or reliability in different conditions. Moreover, there are different underlying reasons for uncertainty. Uncertainty may be due to inherent unpredictability in a population with the current feature set (aleatoric or irreducible uncertainty) or instead due to limited training data (epistemic or reducible uncertainty). Distinguishing these types of uncertainty is important in order to take the correct downstream action. High aleatoric uncertainty indicates that the current prediction task is inherently challenging, and we cannot have high confidence in any one prediction unless we are able to collect richer data (not just more data). On the other hand, high epistemic uncertainty signals that collecting more training data in this region is likely to improve the model, which makes it suitable for guiding model development and active learning (Houlsby et al., 2011; Gal et al., 2017; Smith et al., 2023). We focus on this quantity in this work.
@@
A popular measure of epistemic uncertainty is the mutual information (MI) between model parameters and predictions which quantifies the expected reduction in uncertainty from observing new data (Houlsby et al., 2011; Depeweg et al., 2018). MI is a fundamentally Bayesian quantity that requires access to the posterior distribution over model parameters, which generally renders its exact calculation intractable. As such, substantial effort has gone into developing approximate Bayesian computational techniques (MacKay, 1995; Graves, 2011; Welling and Teh, 2011; Hernandez-Lobato and Adams, 2015; Gal and Ghahramani, 2016), but they are often difficult to scale, compromise predictive accuracy or have weak Bayesian justification.
In light of these challenges, practitioners often use deep ensembles (Lakshminarayanan et al., 2017) to obtain uncertainty estimates. Deep ensembles proceed by generating an ensemble of models by only varying the randomization in the training procedure (e.g. initial weights, data shuffling etc.). Previous works have noted that probabilistic predictions from the deep ensemble reliably quantify uncertainty and exhibit strong performance on downstream tasks (Lakshminarayanan et al., 2017; Ovadia et al., 2019; Ashukha et al., 2020). While Bayesian interpretations have been proposed (Wilson and Izmailov, 2020), a theoretical frequentist understanding behind the success of deep ensembles is lacking. Furthermore, various sources such as limited data, misspecified models or noisy optimization contribute to the uncertainty (Hüllermeier and Waegeman, 2021) and estimates may only account for some of these. Therefore, to correctly use an uncertainty estimate, it is crucial to understand the sources it captures. Our work aims to provide a frequentist perspective on the nature of epistemic uncertainty captured by deep ensembles.
Our analysis proceeds in three parts. Firstly, we provide an asymptotic expansion for MI which links it to the Fisher information and shows how epistemic uncertainty differs from frequentist variance. We use this to motivate the construction of a bootstrap estimator which targets the MI. Secondly, since the epistemic uncertainty within a deep ensemble arises only due to the randomness in the training procedure, we decompose the bootstrap estimator into components arising due to data sampling and training randomness to facilitate comparison. Finally, through empirical studies, we note that deep ensembles capture the portion of epistemic uncertainty due to stochastic optimization, which in turn appears to be the majority component, explaining their success.
@@ 1.1 Our Contribution
Our contributions can be summarized as follows.
@@
  • Contributions to MI from data variability and training randomness: We use the asymptotic expansion to construct a bootstrap estimator of MI. To allow comparison with deep ensembles, we show how this bootstrap estimator separates into components arising from training stochasticity and variability in data sampling.
  • Empirical comparison with deep ensembles: We conduct empirical studies across a variety of deep learning tasks that show that training stochasticity is dominant over variability in data sampling, with deep ensembles capturing the latter.
@@ 1.2 Illustration
To provide the key intuition motivating the bootstrap estimator, we provide the following simulation. Consider a binary logistic regression model, $\mathbb{P}(y=1|x,\theta)=1/(1+e^{-\theta_{1}x-\theta_{0}})$. We sample the features $x$ from a standard normal and also impose standard normal priors on $\theta_{0},\theta_{1}$. Fig. 1(left) visualizes the data and the true model along with 100 posterior samples from the true Bayesian posterior, obtained using a Markov chain Monte Carlo (MCMC) algorithm. Next, we calculate maximum likelihood estimates (MLEs) for the parameters across 100 bootstrapped datasets and show these in Fig. 1(middle)—note the remarkable similarity in the posterior and bootstrap samples. Finally, we use these MCMC and bootstrap samples to form our estimate of the epistemic uncertainty as per Algorithm 1 noting the excellent agreement in Fig. 1(right). We leverage this connection to study Bayesian MI using bootstrapping, which then allows a comparison with deep ensembles.
@@ 2.1 Mutual Information as Epistemic Uncertainty
A popular uncertainty decomposition (MacKay, 1992; Houlsby et al., 2011; Depeweg et al., 2018) leverages information-theoretic quantities to propose measures of aleatoric and epistemic uncertainty. Let the random variables $Y$ and $X$ denote the labels and features respectively in a supervised learning setup. Let $\mathcal{D}_{n}$ be the data comprising of $n$ realizations of $(Y,X)$. Further, let $\theta$ be the random variable denoting the parameters of a prediction model. Information theory states that $\mathrm{H}(Y)$ is the uncertainty in $Y$ and defines mutual information as
@@
The quantification of epistemic uncertainty as the mutual information (MI) between the (test) data and parameters is conceptually pleasing since Eq. 1 reveals it to be the reduction in label uncertainty from knowing the model parameters. This MI is a fundamentally Bayesian object since it requires $\theta$ to be random.
@@ 2.2 Bootstrap
The original bootstrap procedure (Efron, 1979) was proposed to estimate the variance of an estimator under different draws of the data. Consider a setting where there are no repeated data points. Bootstrap first samples weights from a symmetric multinomial distribution with all event probabilities $1/n$. These weights are used to reweight the original data to create a “bootstrapped dataset” over which the estimator is recomputed. This procedure is replicated $B$ times to obtain $B$ “bootstrapped estimates”. The key intuition behind the bootstrap—the “plug-in” principle—proposes that the distribution of these bootstrapped estimates approximates the distribution of the original estimator under redraws of the data. As such, we can obtain an estimate of the variance of our estimator as the empirical variance of the bootstrapped estimates.
We can also instead draw weights from a symmetric Dirichlet distribution (Rubin, 1981). We hypothesize that in a deep learning setting, these weights are more suitable since it allows the neural network to still see all the training instances and hence capture the training data better than multinomial weights which omit training points entirely. For our paper, we thus use Dirichlet weights for our experiments but emphasize that our theorems hold for multinomial weights too (Appendix A).
@@ 2.3 Deep Ensembles
Deep ensembles (Lakshminarayanan et al., 2017) are a simple but powerful uncertainty quantification technique for deep neural networks (DNNs). A deep ensemble is obtained by training multiple models with the exact same architecture and training data but with different random seeds controlling stochasticity in the training procedure (different parameter initializations, different data shuffling when using stochastic gradient descent etc.). The deep ensemble model prediction is simply the average of predictions of member models. They have shown success at a variety of uncertainty quantification tasks (Lakshminarayanan et al., 2017; Ovadia et al., 2019; Ashukha et al., 2020).
@@ 3.1 Problem Setup
$\mathrm{I}\left(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n}\right)$
$=$ $\mathrm{H}\left(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n}\right)-\mathrm{H}\left(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n}\right)$
$=$ $\mathrm{H}\left(\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\hat{p}(X_{\mathrm{test}};\theta)\right]\right)-\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))\right].$
@@ 3.2 Asymptotics of Mutual Information
Theorem 3.1.
Under the assumptions needed for Bernstein von-Mises to hold and uniform integrability of the predicted probabilities (1-7 in Appendix A) we have the asymptotic expansion
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=\frac{1}{2n}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}+o_{p}(n^{-1})$ (2)
$\forall\ X_{\mathrm{test}}$, where $\sigma_{k}^{2}$ is
@@
Proof Sketch: Our analysis proceeds by Taylor expanding the entropy function about the mean posterior predicted probability. We then apply the functional delta method to the Bernstein von-Mises theorem to obtain the limiting law of our predictions. Recall that the Bernstein von-Mises theorem holds under correct model specification, sufficient smoothness, identifiability, and feasibility of $\theta_{0}$ under the prior (Theorem 10.1 of (van der Vaart, 2000)). Finally, we use uniform integrability to obtain the limiting posterior moments. We defer the full proof to Appendix A.
We briefly digress to comment on the theorem and provide better intuition for epistemic uncertainty as measured by mutual information. Firstly, the theorem shows that mutual information decreases linearly in the number of samples. Next, it connects mutual information to the Fisher information and perhaps surprisingly shows how the first order asymptotics are governed by the inverse Fisher information divided by the true conditional probabilities.
@@
as noted in (MacKay, 1992). Recall that $n^{-1}\mathcal{I}^{-1}(\theta_{0})$ is the (asymptotic) variance of the MLE. The numerator in the expression above can then be interpreted as the variance of the model prediction at $X_{\mathrm{test}}$. 1/2p(1p)1/2p(1-p) achieves a minimum at $1/2$ and tends to infinity at 0 and $1$. The first order term is then directly proportional to the model variance but also inversely proportional to the the ground truth certainty (i.e. how close $\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})$ is to 0 or 1). Thus, we see that this first order term matches the intuition behind epistemic uncertainty; if our predictions are uncertain but the true outcome is certain then we can expect to reduce our uncertainty and hence have high epistemic uncertainty.
3.3 A Bootstrap Estimator
Theorem 3.1 suggests that asymptotically correct estimation of the Fisher information and true model probabilities can thus be used to construct first order accurate estimators of mutual information. In light of this, we consider a bootstrap based estimator which we denote by $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$.
Estimation proceeds by first sampling weights $\xi$ from the symmetric Dirichlet with concentration parameter 1. Denote the MLE of $\theta$ over the bootstrapped dataset as $\hat{\theta}_{b}$ (for example, as obtained by minimizing cross-entropy loss). We denote the distribution of this estimate (induced by the randomness in the weights) given the data $\mathcal{D}_{n}$ as $p(\hat{\theta}_{b}|\mathcal{D}_{n})$. Then,
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n}):=$
$\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
Algorithm 1 makes this procedure explicit for clarity.
@@ 3.3 A Bootstrap Estimator
Theorem 3.2.
Under the assumptions needed for Bernstein von-Mises to hold, for asymptotic normality of the MLE to hold and uniform integrability of the predicted probabilities (1-12 in Appendix A), we have
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}^{n}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}}.$
Proof Sketch: We proceed as in Theorem 3.1 but invoke bootstrap limit laws (Theorem 10.16 of (Kosorok, 2007)) instead of Bernstein von-Mises to obtain an identical asymptotic expansion for our estimator. We provide the complete proof in Appendix A.
@@
Theorem 3.2 proves the asymptotic validity of the bootstrap estimator and provides a frequentist path to studying the Bayesian MI.